OpenStudy (anonymous):

Can someone help me with Cal 2? int. by parts

OpenStudy (anonymous):

you got a specific question?

OpenStudy (anonymous):

I have two examples

OpenStudy (anonymous):

general idea is \[\int udv=uv-\int vdu\]

OpenStudy (anonymous):

\[\int\limits x(x+7)^{30}dx\]

OpenStudy (anonymous):

i think ( i suck at these) that you do not use parts for this, but rather a u - sub lets try that

OpenStudy (anonymous):

put \(u=x+7, du = dx, x=u-7\) and turn it in to \[\int (u-7)u^{30}dyu=\int u^{31}du -\int 7u^{30}du\]

OpenStudy (anonymous):

typo there \[\int (u-7)u^{30}du=\int u^{31}du -\int 7u^{30}du\]

OpenStudy (anonymous):

my u=x and my v1= (x+7)30

OpenStudy (anonymous):

trying to figure out my u1= and v=

OpenStudy (anonymous):

i really don't think that will make it any easier use the u -sub and you should be fine

OpenStudy (anonymous):

lets try it \[\int udv=uv-\int vdu\] \[u=x, du = dx, dv =(x+7)^{30}, v = \frac{1}{31}(x+7)^{31}\]

OpenStudy (anonymous):

then you get \[\frac{1}{31}x(x+7)^{31}-\frac{1}{31}\int (x+7)^{31}dx\]

OpenStudy (anonymous):

that might actually work, i see what you would get the same answer as the one i would get above

OpenStudy (anonymous):

where does the 31 come from??

OpenStudy (saifoo.khan):

You have to increase the exponent by 1 and divide it by th same number, so 30+1 = 31.

OpenStudy (anonymous):

Okay so can you show me a step by step to get the final answer? I am a little lost.

zepdrix (zepdrix):

Satellites steps were pretty clear, were you confused by them? I'll go into a little bit more detail if you need. \[\large \int\limits \color{green}{x}(\color{royalblue}{x+7})^{30}\color{purple}{dx}\] Make a u-substitution,\[\large \color{royalblue}{u=x+7}\]Within our substitution here, let's subtract 7 from each side,\[\large \color{green}{x=u-7}\] We also want to take the derivative of the blue term,\[\large \color{purple}{du=dx}\] Plugging all the pieces in gives us,\[\large \int\limits\limits \color{green}{x}(\color{royalblue}{x+7})^{30}\color{purple}{dx} \qquad=\qquad \large \int\limits\limits \color{green}{(u-7)}(\color{royalblue}{u})^{30}\color{purple}{du}\] Make sense so far? This method is U-substitution. If you prefer Integration By Parts, let me know so I don't waste my time with this method XD

OpenStudy (anonymous):

ok it starting to make more sense with the different colors

zepdrix (zepdrix):

From here, we distribute the u^30 to each term in the first set of brackets.\[\large \int\limits u\cdot u^{30}-7u^{30}\;du\]So this is what we're left integrating.\[\large \int\limits u^{31}-7u^{30}\;du\] From here we just apply the `Power Rule for Integration`. Remember how to do that?

OpenStudy (anonymous):

now i see how you get 31

OpenStudy (anonymous):

you add the U's

zepdrix (zepdrix):

Yah it's a rule of exponents that is good to remember! :)\[\large x^2\cdot x^3=x^{2+3}\]

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

our next step would be?

zepdrix (zepdrix):

Apply the Power Rule for Integration to each term. Example:\[\large \int\limits x^2\;dx \qquad=\qquad \frac{x^{2+1}}{2+1} \qquad=\qquad \frac{1}{3}x^3\] 2 Steps: ~Increase exponent by 1. ~Divide by this new exponent.

OpenStudy (saifoo.khan):

@zepdrix the boss!

zepdrix (zepdrix):

:O

OpenStudy (saifoo.khan):

O:

zepdrix (zepdrix):

Simmer down Genghis XD

OpenStudy (anonymous):

ok :)

zepdrix (zepdrix):

Did you try it? :D What'd you get?

OpenStudy (anonymous):

1/31

OpenStudy (saifoo.khan):

:(

zepdrix (zepdrix):

wut..?

zepdrix (zepdrix):

1/31? You should get more than just that :o Where are you u's and stuff?

OpenStudy (saifoo.khan):

Lol. Nothing.

OpenStudy (anonymous):

\[(u ^{30}-7^{30}) (u)^{30}\]

OpenStudy (anonymous):

and 1 for the u outside the bracket \[u(u ^{30}-7^{30})(u)^{30}\]

zepdrix (zepdrix):

I can't figure out what you're doing... hmm

zepdrix (zepdrix):

If we look at just the first term we had,\[\large \int\limits u^{31}du\]We increase the exponent by 1, then divide by this new exponent.\[\large =\qquad\frac{u^{31+1}}{31+1} \qquad=\qquad\frac{1}{32}u^{32}\]

OpenStudy (anonymous):

okay then what is next

zepdrix (zepdrix):

Try to do the same for the next term.\[\large \int\limits 7u^{30}du\]The 7 is a constant, let's pull it out of the integral and ignore it for now.\[\large 7\int\limits u^{30}du\]So applying the rule above, what do you get for this one? :o

OpenStudy (anonymous):

\[\frac{ 1 }{ 31 }u ^{31}\]

zepdrix (zepdrix):

k good. So our integral gives us,\[\large \int\limits u^{31}-7u^{30}\;du \qquad=\qquad \frac{1}{32}u^{32}-\frac{7}{31}u^{31}+C\]

zepdrix (zepdrix):

Then as a final step, undo your substitution. Replace your u's with (x+7)'s.

OpenStudy (anonymous):

so to get 32 you do that step over again?

zepdrix (zepdrix):

yes, apply that rule to the first term gave us the 32 power and fraction.

OpenStudy (anonymous):

i have one more problem can you help me?